# Mass-loaded magnetic explosions in the context of Magnetar Giant Flares and Fast Radio Bursts

### v1

@@ 6 Conclusions
We have presented a class of self-similar, axisymmetric MHD solutions describing relativistic, magnetically dominated explosions with mass loading, motivated by energetic transient phenomena in magnetar magnetospheres. The solutions encompass a broad range of behaviours, characterised by different degrees of baryon loading and pressure support, while remaining analytically tractable.
@@
A central result of this work is that the early-time evolution of these explosions is generically dominated by electromagnetic energy, with kinetic and thermal components becoming important only as a consequence of mass loading and interaction with the external medium. This regime naturally produces rapid, high-energy emission on millisecond timescales, followed by a longer decay associated with magnetic energy depletion and eventual deceleration by a compression shock. The resulting light curves are strongly asymmetric, with rise times controlled primarily by the expansion speed and geometry rather than by the total energy release.
@@
Applied to magnetar giant flares, the model reproduces key observational properties, including sub-millisecond rise times, hard prompt spectra, low radiative efficiencies, and the presence or absence of afterglow emission depending on the ejecta mass. Heavily mass-loaded solutions are associated with efficient energy transfer to shocks and long-lived radio nebulae, while lightly loaded solutions correspond to “clean” magnetospheric reconfigurations with weak or undetectable afterglows.
@@
In this context, FRBs arise naturally from low-mass, highly magnetized explosions in which coherent radio emission can escape before significant baryon loading or shock formation suppresses it. The model therefore predicts that FRBs need not accompany the most energetic magnetar flares, but instead preferentially occur in events where the electromagnetic energy remains dominant throughout the early expansion. This provides a unified explanation for the observed diversity of magnetar activity, including radio-quiet giant flares, radio-loud but energetically modest bursts, and rare events exhibiting both high-energy and radio emission.
@@
The present model is based on a semi-analytical approach, considering ideal MHD, axial symmetry and self-similarity. These simplifications allow the consistent formation of models and the thorough exploration of the parameter space that illustrates the main qualitative characteristics. However, some of the key effects that are associated to observable properties of magnetar flares and FRBs require the relaxation of these idealisations. Namely, the expansion of a spherical magnetised plasmoid, as it propagates away from the neutron star, is not guaranteed to remain self-similar. Moreover, these spherical large-scale plasmoids may deviate from the spherical shape and axial symmetry. Finally, resistive instabilities and the possible formation of shocks due to the interaction of the mass with the external medium will shape the radiative properties of the plasmoid. These deviations can be traced with a numerical simulation that will allow to determine the limitations of the approach made here and their consequences for the observables.
@@
Overall, the results presented here demonstrate that a single, physically motivated class of magnetically dominated, mass-loaded explosions can account for a wide range of observed transient phenomena associated with magnetars. By explicitly linking the dynamics to a small number of physical parameters, the model provides a coherent framework for interpreting present observations and for guiding future multi-wavelength and polarization measurements of magnetar-driven transients.

### v2

@@ 6 Conclusions
We have presented a class of self-similar axisymmetric MHD solutions describing relativistic magnetically dominated explosions with mass loading, motivated by energetic transient phenomena in magnetar magnetospheres. The solutions encompass a broad range of behaviours, characterised by different degrees of baryon loading and pressure support, while remaining analytically tractable.
@@
A central result of this work is that the early-time evolution of these explosions is generically dominated by electromagnetic energy, with kinetic and thermal components becoming important only as a consequence of mass loading and interaction with the external medium. This regime naturally produces rapid high-energy emission on millisecond timescales, followed by a longer decay associated with magnetic energy depletion and eventual deceleration by a compression shock. The resulting light curves are strongly asymmetric, with rise times controlled primarily by the expansion speed and geometry and not by the total energy release.
@@
Applied to magnetar giant flares, the model reproduces key observational properties, including sub-millisecond rise times, hard prompt spectra, low radiative efficiencies, and the presence or absence of afterglow emission, depending on the ejecta mass. Heavily mass-loaded solutions are associated with efficient energy transfer to shocks and long-lived radio nebulae, while lightly loaded solutions correspond to clean magnetospheric reconfigurations with weak or undetectable afterglows.
@@
In this context, FRBs arise naturally from low-mass highly magnetised explosions in which coherent radio emission can escape before significant baryon loading or shock formation suppresses it. The model therefore predicts that FRBs need not accompany the most energetic magnetar flares, but instead preferentially occur in events where the electromagnetic energy remains dominant throughout the early expansion. This provides a unified explanation for the observed diversity of magnetar activity, including radio-quiet giant flares, radio-loud but energetically modest bursts, and rare events exhibiting both high-energy and radio emission.
The present model is based on a semi-analytical approach, considering ideal MHD, axial symmetry, and self-similarity. These simplifications allow the consistent formation of models and the thorough exploration of the parameter space that illustrates the main qualitative characteristics. The qualitative behaviour of the solutions we presented is consistent with the numerical simulations of magnetically driven explosions in mass-loaded environments by Barkov et al. (2022). In particular, both approaches show that increasing the toroidal-to-poloidal magnetic field ratio leads to more compact configurations and a reduction in the maximum expansion velocity. Our semi-analytical solutions complement these numerical results by providing self-consistent equilibrium plasmoid structures that explicitly relate magnetic twist, plasma pressure, and inertia to the global dynamics of relativistic magnetic explosions. However, as these solutions are force-free while in our solutions, pressure and mass are central, a direct comparison is not yet possible.
@@
The key effects that are associated with observable properties of magnetar flares and FRBs require the relaxation of the assumed idealisations we used here. Namely, the expansion of a spherical magnetised plasmoid, as it propagates away from the neutron star, is not guaranteed to remain self-similar. Moreover, these spherical large-scale plasmoids can deviate from the spherical shape and axial symmetry. Finally, resistive instabilities and the possible formation of shocks due to the interaction of the mass with the external medium shape the radiative properties of the plasmoid. These deviations can be traced with a numerical simulation that will allow us to determine the limitations of the approach made here and their consequences for the observables.
@@
Overall, the results presented here demonstrate that a single physically motivated class of magnetically dominated mass-loaded explosions can account for a wide range of observed transient phenomena associated with magnetars. By explicitly linking the dynamics to a small number of physical parameters, the model provides a coherent framework for interpreting current observations and for guiding future multi-wavelength and polarisation measurements of magnetar-driven transients.